OpenStudy (anonymous):

1. Which body is in equilibrium? a. A car racing around a circular track b. A ball falling freely toward the surface of Earth. c. A bike moving with a constant speed along a straight, level road. d. A rocket at the highest point in its trajectory 2. A 50-kilogram roller skate exerts a 20.-newton force on a 80 kilogram roller skater for 0.95 seconds. What is the magnitude of the impulse applied to the 80-kilogram roller skater? 3. When a light ray passes from water (n=1.333) into diamond (n-2.419) at an angle of 45degrees, its path gets bent towards the normal, or away from the n

OpenStudy (anonymous):

1). I believe it would be D. If not D, I think C.

OpenStudy (anonymous):

It's C, since the velocity is constant there is no accelaration on bike, so F=m*a=0,the object could me moving but still in equilibrium.

OpenStudy (anonymous):

I=F*t

OpenStudy (anonymous):

imagine the ligth as a roller car

OpenStudy (anonymous):



OpenStudy (anonymous):

when weel 1, reach the sandy his velocity will decrease, and the weel 2 will be in same velocity as before, so the car will rotate toward the normal, the sandy has a higher refractive index.

OpenStudy (missmob):

for question one answer is b because of terminal velocity

OpenStudy (anonymous):

@missMob a falling body is not in equilibrium since it has an accelaration, and then a resultant force, as stated in second newton's law.

OpenStudy (anonymous):

constant velocity = equilibrium